# Entanglement Microscopy: Tomography and Entanglement Measures via Quantum Monte Carlo

### 22 Feb 2024 v1

@@ V Discussion
Note added.— During the completion of this manuscript, we became aware of a work that shared a similar idea on sampling RDM in spin models [67].
Acknowledgments.—We acknowledge discussions with Gilles Parez. T.T.W., M.H.S. and Z.Y.M. acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region (SAR) of China (Project Nos. 17301721, AoE/P701/20, 17309822, C7037-22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A HKU703/22) and the HKU Seed Funding for Strategic Interdisciplinary Research. W.W.-K. and L.L. are supported by a grant from the Fondation Courtois, a Chair of the Institut Courtois, a Discovery Grant from NSERC, and a Canada Research Chair. We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper.
@@
In this supplementary material, we will provide some additional data on 1d and 2d TFIM in Sec. I, and 2d fermion $t-V$ model in Sec. II.
In Sec. I, we will show the 1d result on the LN $\mathcal{E}^{b}$ and systematic comparison with ED in addition to the 2d result, which has been covered in the main text. We have also done a finite size convergence test in several parameter settings.
In Sec. II, we will show some observables we collected in the $t-V$ model with $L=\beta$. Finite-size scaling, single-particle gap scaling and Green’s function decay are done to justify the location of QCP. LN in this setting is also computed and shown. Then, we will show the finite-temperature data including a similar convergence test and finite size scaling to locate the finite-temperature CPs. Lastly, a color code plot of $\mathcal{E}^{b}$ is presented.
I 1d & 2d Transverse Field Ising Model
@@
I.1 Comparison with Exact Diagonalisation
@@ I.1.1 $T=0$
Fig. S3 (b) shows the two-site $\mathcal{E}^{b}$ as a function of the transverse field $h$. At both distance $r=1$ and $r=\sqrt{2}$, $\mathcal{E}^{b}$ peak at around $h=3$, close to the QCP. At large $h$, $\mathcal{E}^{b}(1)\sim 1/h^{0.9}$ and $\mathcal{E}^{b}(\sqrt{2})\sim 1/h^{4.6}$. For small $h$, $\mathcal{E}^{b}(\sqrt{2})$ has a sudden death at $h\sim 0.5$.
@@ I.1.2 Finite T
Fig. S3 (c) shows the decay of $\mathcal{E}^{b}(1)$ with temperature, which is very similar to the $L=40$ QMC result in Fig. S3 (c), in terms of the shape of the three curves and the magnitude. There is a small difference in the sudden death temperature, likely due to the finite size effect.
I.2 Example of RDM
Here we give an example result of 3-site RDM for the 2d TFIM with $L=\beta=24$ and $h=3$. The three sites are chosen the same way as we did in investigating tripartite entanglement, and form on the smallest right triangle with the first site on the right angle.
@@ I.2 Example of RDM
II Fermion $t-V$ model
II.1 The $N=4$ GNY quantum critical point
@@
Figure S8: Nearest neighbor LN $\mathcal{E}^{f}$ against $V$. In all cases, inverse temperature $\beta$ is set to equal to the system size. The red star in panel is the ED result of free fermion model with system size and inverse temperature $L=\beta=1000$. Panel (b) is a zoom-in plot of the square region in (a). Panels (a) and (b) share the same legend in (a)
@@
II.2 $\mathcal{E}^{f}$ in finite temperature
@@
Fig. S9 shows the nearest neighbor LN $\mathcal{E}^{f}$ at finite temperature, with $V=0.20$ (DSM phase), $V=1.28$ (near QCP), and $V=3.00$ (CDW phase) respectively. The result of high-temperature fitting yields powers close to 2, which is the theoretical prediction in free case by Ref [32]. a
@@
If one further separate the two subregions, the power will increase as $2r$ as shown in the fitting in Fig. S10. In high temperature, the fermionic Green’s function expands as
$G=1-\frac{1}{1+e^{-\beta H}}=\frac{1}{2}-\frac{\beta}{4}H+\frac{\beta^{3}}{48}H^{3}-\frac{\beta^{5}}{480}H^{5}+\ldots,$
and the lowest order attained by any $\langle c^{\dagger}_{i}c_{j}\rangle$ depends on the minimum number of hopping terms connecting them, which is simply the bond-distance $r_{ij}$. That is, in cases like ours where only nearest hopping exists, one has
$\langle c^{\dagger}_{i}c_{j}\rangle\sim\beta^{r_{ij}}=1/T^{r_{ij}}$
In high temperature, $\mathcal{E}^{f}$ goes as $1/G_{ij}^{2}\sim 1/T^{2r_{ij}}$ as confirmed by fitting in Fig. S9 and S10

### 18 Mar 2024 v2

@@ Discussion
Methods
Stochastic series expansion.
In the usual SSE simulation [29], the path-integral is Taylor expanded as
$\begin{split}Z&=\operatorname{Tr}\{e^{-\beta H}\}=\sum_{\psi}\bra{\psi}e^{-\beta H}\ket{\psi}\\ &=\sum_{\{b_{1},b_{2}\cdots b_{M}\}}\frac{\beta^{n}(M-n)!}{M!}\sum_{\psi}\bra{\psi}H_{b_{M}}\cdots H_{b_{2}}H_{b_{1}}\ket{\psi},\end{split}$ (7)
where $\{b_{1},b_{2}\cdots b_{M}\}$ denotes the type of operators on $M$ imaginary time slices. The sampling of bonds and states is carried out by performing diagonal update, which add and remove diagonal operators, and loop update, which interchange diagonal and off-diagonal operators.
In sampling RDM in SSE, where imaginary time boundary condition is opened for sites in region $A$, one simply break the loop at these points in loop update, and the detailed balance condition will be satisfied. Hence, the frequency of any edge configuration is proportional to the corresponding element in the RDM.
Determinant quantum Monte Carlo. For the $t-V$ model, we employ the determinant quantum Monte Carlo (DQMC) scheme [30, 31, 32], where the fermion interaction is decoupled into fermion bilinear coupled to auxiliary field. In each configuration, the free fermionic Green’s function can be calculated, and higher order correlation is evaluated using Wick’s theorem.
Acknowledgments
We acknowledge discussions with Gilles Parez. T.-T.W., M.H.S. and Z.Y.M. acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region (SAR) of China (Project Nos. 17301721, AoE/P701/20, 17309822, C7037-22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A HKU703/22) and the HKU Seed Funding for Strategic Interdisciplinary Research. W.W.-K. and L.L. are supported by a grant from the Fondation Courtois, a Chair of the Institut Courtois, a Discovery Grant from NSERC, and a Canada Research Chair. We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper.
@@ Acknowledgments
In this supplementary material, we will provide some additional data on 1d and 2d TFIM in Sec. 1d & 2d Transverse Field Ising Model, and 2d fermion $t-V$ model in Sec. Fermion $t-V$ model.
In Sec. 1d & 2d Transverse Field Ising Model, we will show the 1d result on the LN $\mathcal{E}^{b}$ and systematic comparison with ED in addition to the 2d result, which has been covered in the main text. We have also done a finite size convergence test in several parameter settings.
In Sec. Fermion $t-V$ model, we will show some observables we collected in the $t-V$ model with $L=\beta$. Finite-size scaling, single-particle gap scaling and Green’s function decay are done to justify the location of QCP. LN in this setting is also computed and shown. Then, we will show the finite-temperature data including a similar convergence test and finite size scaling to locate the finite-temperature CPs. Lastly, a color code plot of $\mathcal{E}^{b}$ is presented.
1d & 2d Transverse Field Ising Model
@@
Comparison with Exact Diagonalisation
@@
Fig. S3 (b) shows the two-site $\mathcal{E}^{b}$ as a function of the transverse field $h$. At both distance $r=1$ and $r=\sqrt{2}$, $\mathcal{E}^{b}$ peak at around $h=3$, close to the QCP. At large $h$, $\mathcal{E}^{b}(r)$ decays algebraically for $r=1$ and $\sqrt{2}$.
@@
Fig. S3 (c) shows the decay of $\mathcal{E}^{b}(1)$ with temperature, which is very similar to the $L=40$ QMC result in Fig. S3 (c), in terms of the shape of the three curves and the magnitude. The sudden death temperatures are different from that of QMC, likely due to the finite size effect.
Example of RDM
Here we give an example result of 3-site RDM for the 2d TFIM with $L=\beta=24$ and $h=3$. The three sites are chosen the same way as we did in investigating tripartite entanglement, and form the smallest right triangle with the first site on the right angle. The RDM reads:
@@
Fermion $t-V$ model
The $N=4$ GNY quantum critical point
@@
Figure S8: Nearest neighbor LN $\mathcal{E}^{f}$ against $V$. In all cases, inverse temperature $\beta$ is set to equal to the system size. The red star in panel is the exact result of the free fermion model with system size and inverse temperature $L=\beta=1000$. Panel (b) is a zoom-in plot of the square region in (a). Panels (a) and (b) share the same legend in (a)
@@
$\mathcal{E}^{f}$ in finite temperature
@@
Fig. S9 shows the nearest neighbor LN $\mathcal{E}^{f}$ at finite temperature, with $V=0.20$ (DSM phase), $V=1.28$ (near QCP), and $V=3.00$ (CDW phase) respectively. The result of high-temperature fitting yields powers close to 2, which is the theoretical prediction in the free case [37].
@@
If one further separate the two subregions, the power will increase as $2r$ as shown in the fitting in Fig. S10, see Main text for the explanation.